OpenStudy (anonymous):

x-6÷3=2x+4÷4 what is it please

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

x-2=2x+1 so x= -3

OpenStudy (anonymous):

\[-1-2=2x-x hence x=-3